OpenStudy (anonymous):

An automobile engine slows down from 3500rpm to 1200rpm in 2.5s. calculate (a) its angular acceleration, assumed constant, and (b) the total number of revolutions the engine makes in this time.

OpenStudy (anonymous):

I already did part a, and got -96.34 rad/s^2. However, I am stuck in the second part. I thought i could use this formula but I keep getting it wrong. \[\theta= \omega _{i}t+ \frac{ 1 }{ 2 } \alpha t^{2}\] then convert it to revolutions by dividing theta by 2pi.